# QuadGPT: Native Quadrilateral Mesh Generation with Autoregressive Models

### v1

@@ 1 Introduction
In the pursuit of automated generation of artist-ready 3D assets from inputs such as text or images, existing approaches typically decouple the problem into two separate stages: geometry generation and topology generation. For geometry, latent diffusion models (LDMs) have achieved remarkable success in generating 3D shapes (Zhao et al., 2025b; Li et al., 2025d) via implicit representations like SDF (Li et al., 2024). Meshes are then extracted with iso-surface algorithms (Lorensen & Cline, 1998; Shen et al., 2021; 2023), inevitably resulting in unstructured dense meshes. On the other hand, topology-focused methods using cross-field guidance (Huang et al., 2018; Dong et al., 2025a) aim for structure but are often not robust, requiring pristine input meshes and failing to produce the adaptive, artist-like tessellation where polygon density matches geometric complexity.
@@ 2 Related Work
A dominant paradigm in 3D generation relies on continuous neural representations like vecsets (Zhang et al., 2023; 2024; Li et al., 2024; Zhao et al., 2023; Wu et al., 2024; Li et al., 2025b; Zhao et al., 2025b) or sparse voxels (Xiang et al., 2024; Ye et al., 2025; He et al., 2025; Wu et al., 2025; Li et al., 2025d). A universal limitation of these methods is their reliance on an iso-surfacing step, such as Marching Cubes (Lorensen & Cline, 1998), which invariably yields dense, topologically unstructured triangular meshes. Concurrently, traditional approaches to quadrilateral meshing are predominantly guided by cross-field computation. These methods are either optimization-based, requiring slow, per-shape optimization (Bommes et al., 2009; Jakob et al., 2015; Huang et al., 2018; Bommes et al., 2013; Ebke et al., 2016; Campen & Kobbelt, 2014; Jiang et al., 2015; Diamanti et al., 2015), or more recent learning-based techniques that accelerate field prediction (Dielen et al., 2021; Li et al., 2025c; Dong et al., 2025b; a). However, all field-guided methods depend on multi-stage pipelines that are not end-to-end generative frameworks.
@@ 3.2 Autoregressive Pre-training
Rather than treating mesh generation as a generic sequence task, we leverage the inherent hierarchical structure of mesh data. To this end, we utilize the Hourglass Transformer architecture (Hao et al., 2024; Nawrot et al., 2021), which processes the input sequence at multiple levels of abstraction. Let the input token sequence embeddings be $\mathbf{E}^{(0)}\in\mathbb{R}^{L\times D_{0}}$. The architecture employs a series of causality-preserving shortening layers to create a computational bottleneck. The sequence is first processed by a Transformer block and then shortened by a factor of 4, and subsequently by a factor of 3:
$\mathbf{E}^{(1)}$ $=\mathtt{Shorten}_{4}(\mathtt{TransformerBlock}_{1}(\mathbf{E}^{(0)}))\in\mathbb{R}^{(L/4)\times D_{1}}$ (3)
$\mathbf{E}^{(2)}$ $=\mathtt{Shorten}_{3}(\mathtt{TransformerBlock}_{2}(\mathbf{E}^{(1)}))\in\mathbb{R}^{(L/12)\times D_{2}}$ (4)
@@
Our training strategy combines truncated sequence training for efficiency with a novel curriculum for stability. To manage the long sequences of high-resolution meshes, we employ truncated training, using fixed-length segments (e.g., 36,864 tokens) to enable efficient, large-batch processing. Furthermore, to ensure stable learning of complex quadrilateral topology, we introduce a curriculum learning strategy. We first initialize QuadGPT with weights from a model pre-trained exclusively on triangular meshes. We then progressively fine-tune this model, using our quad-dominance condition $r$ to gradually anneal the training data distribution from purely triangular ($r=0$) to quad-dominant ($r\to 1$). This graduated exposure allows the model to master basic geometric syntax before tackling the more complex rules of quadrilateral topology, significantly improving stability and convergence speed.
@@ 3.3 Topological Refinement with Reinforcement Learning
We fine-tune the pretrained QuadGPT policy by forming preference pairs using our topological rewards and optimizing a DPO-style objective on the collected pairs. Let $\pi_{\theta}$ denote the trainable policy being fine-tuned, $\pi_{\text{ref}}$ the frozen reference model, and $\beta>0$ a parameter controlling the deviation from $\pi_{\text{ref}}$. The current policy $\pi_{\theta}$ produces candidates for input $x$, which are ranked by topological rewards to yield preference pairs $\bigl(y_{w},y_{l}\bigr)$ with
@@ 4.1 Experimental Settings
We pretrain QuadGPT on a cluster of 64 NVIDIA A100 GPUs for 7 days using the AdamW optimizer (Loshchilov & Hutter, 2019) ($\beta_{1}=0.9,\beta_{2}=0.95$) with a learning rate of 1e-4 and a linear warmup schedule. The Decoder is a 1.1B parameter model featuring 24 Transformer layers arranged in a three-stage hourglass architecture. This architecture employs linear downsample layers with factors of 4 and 3 to efficiently process long sequences. The subsequent RL fine-tuning stage is performed for 4 hours on the same hardware setup, using a reduced learning rate of 1e-7. Both pre-training and fine-tuning leverage a truncated sequence strategy to manage the long contexts of high-resolution meshes. During inference, the architecture supports a context window of 36,864 tokens. We generate meshes using a combination of top-k and nucleus (top-p) sampling with $k=10$ and $p=0.95$, along with a temperature of $T=0.5$, to balance output diversity and stability. To optimize performance, inference is accelerated through a custom implementation of KV caching and CUDA graphs, specifically tailored for the hourglass architecture. This achieves a generation speed of approximately 230 tokens per second on a single A100 GPU.
@@
We evaluate QuadGPT against two primary categories of state-of-the-art mesh generation methods. The first category comprises leading autoregressive models that generate triangular meshes, including MeshAnythingV2 (Chen et al., 2025c), BPT (Weng et al., 2025), DeepMesh (Zhao et al., 2025a), and FastMesh (Kim et al., 2025). Since these methods are fundamentally designed to produce triangular outputs, we apply a robust triangle-to-quadrilateral conversion algorithm as a post-processing step to facilitate a fair comparison. The second category represents specialized quad-meshing techniques, for which we include the well-established field-guided method QuadriFlow (Huang et al., 2018). This selection provides a comprehensive benchmark against both the latest in generative modeling and classic, topology-focused approaches.
@@ 4.2 Qualitative Results
We first present a qualitative comparison of QuadGPT against existing baselines across two distinct domains: challenging, out-of-distribution dense meshes generated by other AI models, and in-distribution, high-quality artist-designed meshes. As illustrated in Figure 4, the indirect pipeline of converting triangular meshes from autoregressive baselines often struggles. These methods frequently produce meshes with significant topological artifacts, missing geometric details, or overly simplified structures that fail to capture the nuance of the original shape. In contrast, QuadGPT consistently generates meshes that are significantly more coherent, artistically plausible, and faithful to the input geometry. Our model demonstrates robust performance on the challenging AI-generated assets and achieves near-perfect topological reconstruction on the artist-designed meshes. Figure 5 compares our method against the field-guided approach of QuadriFlow. The baseline exhibits significant instability, particularly on meshes with complex topology, sharp features, or thin structures. This often results in a severe loss of geometric detail or even catastrophic failures. QuadGPT, in contrast, demonstrates exceptional robustness, faithfully reconstructing the geometry while maintaining high-quality, structured topology across all examples.
@@ 4.3 Quantitative Results
Table 1 summarizes the quantitative comparison of QuadGPT against all baselines across both artist-designed and AI-generated dense meshes. We report metrics for geometric fidelity (CD, HD) and topological quality (Quad Ratio, QR), alongside a comprehensive User Study (US) to assess perceptual quality. The results demonstrate that QuadGPT consistently and significantly outperforms competing approaches in both quality and robustness. The user study, where experts ranked the outputs of all six methods from best (5 points) to worst (0 points), reveals a decisive preference for our method. This confirms that its advantages in producing production-ready assets are not only quantitatively measurable but also perceptually significant.
@@ 4.4 Ablation Studies
Figure 6: Effectiveness of tDPO-Pro. Our complete training strategy significantly improves the quality and structural integrity of the generated quad-meshes.
Table 2: Quantitative Evaluation of Training Strategy. The user study, where experts ranked the outputs of all five methods from best (4 points) to worst (0 points)
@@ 4.4.1 Effectiveness of Curriculum-based Pre-training
We validate our curriculum learning strategy by comparing it against training directly on quad-dominant meshes (”From Scratch”). As shown in Table 2, initializing from a converged triangle-generation model (”Finetune”) yields superior geometric fidelity and is strongly preferred by users. This confirms that leveraging a simpler, related task as a warm-up provides a more stable foundation for learning the complex patterns of quadrilateral topology.
4.4.2 Effectiveness of tDPO
We analyze the impact of our DPO components by comparing three variants: DPO (fine-tuning only on low face-count meshes whose full token sequences fit within the context window), tDPO (truncated training with a basic reward penalizing only fractures), and our full tDPO-Pro model (truncated training with the complete topological reward).As shown in Table 2, DPO fails to generalize, confirming the need to fine-tune on complex meshes. tDPO, which leverages our truncated training strategy, dramatically improves performance, especially in the user study. Finally, our tDPO-Pro model achieves the best results across all metrics. This improvement visualized in Figure 6, provides strong empirical support for our complete, topology-aware, truncated DPO framework, demonstrating that the full edge flow reward is crucial for generating high-quality quad-meshes.
@@ 5 Conclusion
In this work, we introduced QuadGPT, the first autoregressive framework capable of directly generating native quadrilateral and mixed-element meshes. We further demonstrated that by pioneering a tDPO fine-tuning stage, our model can explicitly optimize for the global topological quality crucial for production-ready assets. Additionally, we proposed a unified serialization scheme that seamlessly handles mixed-element topologies, enabling a single, scalable model to address the long-standing challenge of structured mesh generation. Our approach establishes a new state-of-the-art in generative meshing, bridging the gap between the power of large-scale sequence models and the stringent topological requirements of professional 3D content creation. This work offers a promising new direction for the automated creation of high-quality, structured 3D assets, paving the way for future research and applications in computer graphics and beyond.
@@ 7 Reproducibility Statement
This paper presents a scalable autoregressive framework for native quad-mesh generation. Although full model release is not feasible at this stage, the method is comprehensively detailed in terms of data representation, network architecture, and training protocols to facilitate replication and extension. We emphasize that high-quality data curation, including rigorous collection, processing, and filtering of open-source data, is essential for achieving comparable performance. To support validation and downstream use, we will provide a public API.
@@ A.3 Final Dataset Compilation
 if num_vertices(component) ¡ $\tau_{\text{vtx\_min}}$ then
@@ Appendix C More Results
Appendix D More Ablation Study

### v2

@@ 1 Introduction
In the pursuit of automated generation of artist-ready 3D assets from inputs such as text or images, existing approaches typically decouple the problem into two separate stages: geometry generation and topology generation. For geometry, latent diffusion models (LDMs) have achieved remarkable success in generating 3D shapes (Zhao et al., 2025b; Li et al., 2025d) via implicit representations like SDF (Li et al., 2024). Meshes are then extracted with iso-surface algorithms (Lorensen and Cline, 1998; Shen et al., 2021; 2023), inevitably resulting in unstructured dense meshes. On the other hand, topology-focused methods using cross-field guidance (Huang et al., 2018; Dong et al., 2025a) aim for structure but are often not robust, requiring pristine input meshes and failing to produce the adaptive, artist-like tessellation where polygon density matches geometric complexity.
@@ 2 Related Work
A dominant paradigm in 3D generation relies on continuous neural representations like vecsets (Zhang et al., 2023; 2024; Li et al., 2024; Zhao et al., 2023; Wu et al., 2024; Li et al., 2025b; Zhao et al., 2025b) or sparse voxels (Xiang et al., 2024; Ye et al., 2025; He et al., 2025; Wu et al., 2025; Li et al., 2025d). A universal limitation of these methods is their reliance on an iso-surfacing step, such as Marching Cubes (Lorensen and Cline, 1998), which invariably yields dense, topologically unstructured triangular meshes. Concurrently, traditional approaches to quadrilateral meshing are predominantly guided by cross-field computation. These methods are either optimization-based, requiring slow, per-shape optimization (Bommes et al., 2009; Jakob et al., 2015; Huang et al., 2018; Bommes et al., 2013; Ebke et al., 2016; Campen and Kobbelt, 2014; Jiang et al., 2015; Diamanti et al., 2015), or more recent learning-based techniques that accelerate field prediction (Dielen et al., 2021; Li et al., 2025c; Dong et al., 2025b; a). However, all field-guided methods depend on multi-stage pipelines that are not end-to-end generative frameworks.
@@ 3.2 Autoregressive Pre-training
Rather than treating mesh generation as a generic sequence task, we leverage the inherent hierarchical structure of mesh data. To this end, we utilize the Hourglass Transformer architecture (Hao et al., 2024; Nawrot et al., 2021), which processes the input sequence at multiple levels of abstraction. Let the input token sequence embeddings be $\mathbf{E}^{(0)}\in\mathbb{R}^{L\times D_{0}}$. The architecture employs a series of causality-preserving shortening layers to create a computational bottleneck. The sequence is first processed by a Transformer block and then shortened by a factor of 3, and subsequently by a factor of 4:
$\mathbf{E}^{(1)}$ $=\mathtt{Shorten}_{3}(\mathtt{TransformerBlock}_{1}(\mathbf{E}^{(0)}))\in\mathbb{R}^{(L/3)\times D_{1}}$ (3)
$\mathbf{E}^{(2)}$ $=\mathtt{Shorten}_{4}(\mathtt{TransformerBlock}_{2}(\mathbf{E}^{(1)}))\in\mathbb{R}^{(L/12)\times D_{2}}$ (4)
@@
Our training strategy combines truncated sequence training for efficiency with a novel curriculum for stability. To manage the long sequences of high-resolution meshes, we employ truncated training, using fixed-length segments (e.g., 36,864 tokens) to enable efficient, large-batch processing. Furthermore, to ensure stable learning of complex quadrilateral topology, we introduce a curriculum learning strategy. We first initialize QuadGPT with weights from a model pre-trained exclusively on triangular meshes. We then progressively finetune this model, using our quad-dominance condition $r$ to gradually anneal the training data distribution from purely triangular ($r=0$) to quad-dominant ($r\to 1$). This graduated exposure allows the model to master basic geometric syntax before tackling the more complex rules of quadrilateral topology, significantly improving stability and convergence speed.
@@ 3.3 Topological Refinement with Reinforcement Learning
We finetune the pretrained QuadGPT policy by forming preference pairs using our topological rewards and optimizing a DPO-style objective on the collected pairs. Let $\pi_{\theta}$ denote the trainable policy being fine-tuned, $\pi_{\text{ref}}$ the frozen reference model, and $\beta>0$ a parameter controlling the deviation from $\pi_{\text{ref}}$. The current policy $\pi_{\theta}$ produces candidates for input $x$, which are ranked by topological rewards to yield preference pairs $\bigl(y_{w},y_{l}\bigr)$ with
@@ 4.1 Experimental Settings
We pretrain QuadGPT on a cluster of 64 NVIDIA A100 GPUs for 7 days using the AdamW optimizer (Loshchilov and Hutter, 2019) ($\beta_{1}=0.9,\beta_{2}=0.95$) with a learning rate of 1e-4 and a linear warmup schedule. The Decoder is a 1.1B parameter model featuring 24 Transformer layers arranged in a three-stage hourglass architecture. This architecture employs linear downsample layers with factors of 4 and 3 to efficiently process long sequences. The subsequent RL fine-tuning stage is performed for 4 hours on the same hardware setup, using a reduced learning rate of 1e-7. Both pre-training and fine-tuning leverage a truncated sequence strategy to manage the long contexts of high-resolution meshes. During inference, the architecture supports a context window of 36,864 tokens. We generate meshes using a combination of top-k and nucleus (top-p) sampling with $k=10$ and $p=0.95$, along with a temperature of $T=0.5$, to balance output diversity and stability. To optimize performance, inference is accelerated through a custom implementation of KV caching and CUDA graphs, specifically tailored for the hourglass architecture. This achieves a generation speed of approximately 230 tokens per second on a single A100 GPU.
@@
We evaluate QuadGPT against two primary categories of state-of-the-art mesh generation methods. The first category comprises leading autoregressive models that generate triangular meshes, including MeshAnythingV2 (Chen et al., 2025c), BPT (Weng et al., 2025), DeepMesh (Zhao et al., 2025a), and FastMesh (Kim et al., 2025). Since these methods are fundamentally designed to produce triangular outputs, we apply a robust triangle-to-quadrilateral conversion algorithm like (Muntoni and Cignoni, 2021) as a post-processing step to facilitate a fair comparison. The second category represents specialized quad-meshing techniques, for which we include the well-established field-guided method QuadriFlow (Huang et al., 2018). This selection provides a comprehensive benchmark against both the latest in generative modeling and classic, topology-focused approaches.
@@ 4.2 Qualitative Results
We first present a qualitative comparison of QuadGPT against existing baselines across two distinct domains: challenging, out-of-distribution dense meshes from other AI models, and in-distribution, high-quality artist-designed meshes. Visual inspection is crucial, as it reveals the subtle yet critical differences in topological quality that metrics alone cannot fully capture. As illustrated in Figure 4, the indirect pipeline of converting triangular meshes from autoregressive baselines often struggles. These methods frequently produce meshes with significant topological artifacts, missing geometric details, or overly simplified structures that fail to capture the original shape’s nuance. In contrast, QuadGPT consistently generates meshes that are significantly more coherent and artistically plausible, producing the clean edge flow characteristic of professional work. Our model demonstrates strong robustness on challenging AI-generated assets and achieves near-perfect topological reconstruction on artist-designed meshes. Figure 5 compares our method against the field-guided approach of QuadriFlow. The baseline exhibits significant instability on meshes with complex topology or sharp features, often resulting in severe geometric degradation or catastrophic failures. QuadGPT, in contrast, demonstrates exceptional robustness, faithfully reconstructing geometry while maintaining high-quality, structured topology in all examples.
@@ 4.3 Quantitative Results
Table 1 summarizes the quantitative comparison of QuadGPT against all baselines across both artist-designed and AI-generated dense meshes. We evaluate geometric fidelity using Chamfer Distance (CD) and Hausdorff Distance (HD), topological quality via Quad Ratio (QR), and perceptual quality through a comprehensive user study (US). The results demonstrate that QuadGPT consistently and significantly outperforms competing approaches in both quality and robustness. The user study, where experts ranked the outputs of all six methods from best (5 points) to worst (0 points), reveals a decisive preference for our method. This confirms that its advantages in producing production-ready assets are not only quantitatively measurable but also perceptually significant.
Figure 6: Effectiveness of tDPO-Pro. Our comprehensive training strategy significantly enhances both the geometric quality and structural integrity of the generated quad-meshes.
Figure 7: Native Generation vs. Conversion Pipeline. (Q) denotes the use of a triangle-to-quad conversion step. Although TriGPT employs the same RL fine-tuning to mitigate fractures, its topological quality is inherently constrained by the post-processing conversion, yielding significantly inferior edge flow compared to our native QuadGPT.
@@ 4.4 Ablation Studies
4.4.1 Effectiveness of tDPO
We analyze the impact of our DPO components by comparing three variants: standard DPO (fine-tuning on full, low-face-count meshes), tDPO (truncated training with a basic fracture penalty), and our full tDPO-Pro model (truncated training with the complete topological reward). Table 2 shows that standard DPO fails to generalize to complex meshes. In contrast, tDPO dramatically improves performance, while tDPO-Pro achieves the best results across all metrics. This improvement, visualized in Figure 6, provides strong empirical support for our comprehensive, topology-aware, truncated DPO framework.
Table 2: Quantitative Evaluation of Training Strategy. User study scores reflect expert rankings of the five methods, ranging from 0 (worst) to 4 (best).
@@ 4.4.2 Effectiveness of Curriculum-based Pre-training
We validate our curriculum learning strategy by comparing it against a model trained directly on quad-dominant meshes from scratch (“From Scratch”). As shown in Table 2, the “From Scratch” model struggles to converge, resulting in poor geometric fidelity. This difficulty arises because predicting a quadrilateral face is inherently more complex than predicting a triangle, as it is topologically equivalent to predicting two correlated triangles simultaneously. Initializing weights from a converged triangle-generation model (“Finetune”) yields significantly superior results. This confirms that our curriculum strategy leveraging the simpler task of triangle generation as a warmup is essential for establishing the stable geometric foundation required to master the more challenging patterns of quadrilateral topology.
4.4.3 Native Generation vs. Conversion Pipeline
To isolate the benefits of our native approach, we introduce a strong baseline, TriGPT, which generates triangles that are then converted to quads. To ensure a fair comparison, TriGPT shares the identical architecture, 1.3 million mesh training dataset, and tDPO reinforcement learning strategy as QuadGPT. This setup controls for all confounding variables, testing only the efficacy of the end-to-end native pipeline versus the generation-then-conversion paradigm.
Table 3: Native vs. Conversion Pipeline. QuadGPT is compared against a strong triangle-generation baseline (TriGPT), both with and without RL. US scores are expert rankings (0–2).
Method CD $\downarrow$ HD $\downarrow$ QR $\uparrow$ US $\uparrow$
TriGPT(Q) 0.062 0.160 70% 0.2
TriGPT+RL(Q) 0.051 0.138 72% 0.5
QuadGPT (Ours) 0.057 0.147 80% 1.3
The results in Table 3 and Figure 7 confirm our hypothesis. While the highly optimized triangle baseline (TriGPT+RL) achieves slightly better geometric scores (CD/HD), it cannot match the topological quality of our native approach. QuadGPT demonstrates a substantially higher Quad Ratio (QR) and, crucially, a user preference score 2.6x higher than the strongest baseline. This decisive gap in perceptual quality confirms that while post-hoc conversion struggles to create coherent global structures, our end-to-end native framework excels at learning the artist-preferred topologies, validating its superior practical utility.
@@ 5 Conclusion
We present QuadGPT, the first autoregressive framework that directly generates native quadrilateral and mixed-element meshes. It achieves state-of-the-art results in generative meshing, setting a new standard for both geometric fidelity and topological quality. Our scalable, neural-first approach departs from previous triangle-based methods and conversion pipelines, which rely on heuristic post-processing to approximate quad topology. By leveraging a unified serialization scheme and a novel topology-aware fine-tuning stage (tDPO), QuadGPT directly optimizes for global structure, making it well-suited for the automated creation of production-ready 3D assets.
@@ 7 Reproducibility Statement
This paper presents a scalable autoregressive framework for native quad-mesh generation. Although full model release is not feasible at this stage, the method is comprehensively detailed in terms of data representation, network architecture, and training protocols to facilitate replication and extension. We emphasize that high-quality data curation, including rigorous collection, processing, and filtering of open-source data, is essential for achieving comparable performance. To support validation and downstream use, we will provide a public API and Code.
Acknowledgments
This research was supported by fundings from the Hong Kong RGC General Research Fund (152228/23E, 162161/24E, 162116/25E, 162180/25E), National Natural Science Foundation of China (NSFC) Key Program (No.62532005), Collaborative Research Fund (No. C1042-23GF, No. 5097-25G), NSFC/RGC Collaborative Research Scheme (Grant No. 62461160332 & CRS_HKUST602/24), Research Impact Fund (No. R5011-23F), Areas of Excellence Scheme (AoE/E-601/22-R), and the InnoHK (HKGAI).
@@ A.3 Final Dataset Compilation
 if $\text{num\_vertices}(\text{component})<\tau_{\text{vtx\_min}}$ then
@@ Appendix C More Results
Appendix D Additional Experimental Analysis
In this section, we provide further analysis to complement the experiments in the main paper.
D.1 Impact of Curriculum Learning and Topology-Aware Rewards
@@
Figure 14: The Irreversibility of Triangulation and Limitations of Conversion Algorithms. (a) A high-quality artist-created quad mesh with clean edge flow. (b) The same mesh after triangulation—the ground truth topological information is now ambiguous. (c-e) The results of applying different tri-to-quad conversion algorithms to (b). While our ILP-based operator (e) produces a better result than the default methods in MeshLab (c) and Blender (d), none can perfectly recover the original topology, resulting in broken edge loops and artifacts. This demonstrates the fundamental limitations of post-hoc conversion.
D.2 The Inherent Limitations of Triangle-to-Quad Conversion
To illustrate the limitations of any post-hoc conversion pipeline, we conduct a best-case scenario experiment. We begin with a high-quality artist-created quad mesh and triangulate it. This process represents an irreversible information loss: the artist’s original topological intent becomes ambiguous, creating a complex combinatorial problem that heuristic algorithms struggle to solve.
As shown in Figure 14, we compare our own ILP-based operator against the default methods in MeshLab Tarini et al. (200) and Blender Blender Foundation (2025). While our optimized operator performs best, none of the methods can perfectly recover the original edge flow, leaving behind topological artifacts.
This challenge is significantly amplified when the input is an AI-generated triangle mesh, which often has an edge flow unsuitable for conversion, as seen in our TriGPT experiments (Figure 7). This compounding effect of a suboptimal input and an imperfect algorithm reinforces the need for our native generation approach.
Furthermore, algorithms from adjacent fields like Blossom-Quad (remacle2011blossom), designed for 2D finite element analysis, are too restrictive for complex 3D surfaces. Its perfect-matching requirement caused it to fail on all but our simplest test cases, underscoring the need for a domain-specific, learning-based solution like QuadGPT.
D.3 Analysis of Compressed vs. Direct Coordinate Representation
A critical design choice in autoregressive mesh generation is the tokenization strategy. Methods like DeepMesh (Zhao et al., 2025a) and BPT (Weng et al., 2025) employ compressed representations. Our experiments with these methods on triangle meshes confirmed a similar finding: compressed tokenizers lead to faster initial convergence due to shorter sequence lengths. However, we argue that faster convergence does not necessarily lead to better final performance. As shown in Table 4, our extensive experiments at the 1B-parameter scale show that a simple, direct coordinate representation ultimately achieves superior results in both geometric fidelity and topological quality, a finding strongly corroborated by our user study. We hypothesize that overly complex, “over-designed” tokenizers can introduce an unintended inductive bias, constraining the model’s expressive capacity and limiting its performance ceiling, especially when trained on large, diverse datasets. This observation aligns with a broader trend in deep learning, demonstrated by models like DepthAnything V3 (Lin et al., 2025), where simple, highly scalable architectures often outperform more complex designs when sufficient data and compute are available. For QuadGPT, we therefore prioritized the higher performance ceiling of the direct representation, confident that techniques like truncated training effectively manage the longer sequence lengths.
Table 4: Quantitative Comparison: Tokenization on Triangle Meshes. Direct coordinate representation (TriGPT) outperforms a compressed BPT-style tokenizer (TriGPT+BPT). User Study scores reflect expert rankings from 0 (worst) to 1 (best). (Q) denotes outputs were converted to quads for evaluation.
Method CD $\downarrow$ HD $\downarrow$ QR $\uparrow$ US $\uparrow$
TriGPT+BPT (Compressed) 0.078 0.198 68% 0.3
TriGPT (Direct Coord.) 0.062 0.160 70% 0.7
D.4 Ablation on Training Data
To disentangle the impact of our methodology from our curated dataset, we conduct a data ablation study. We train a variant, QuadGPT-OS, using only the publicly available portions of our dataset (e.g., ShapeNet Chang et al. (2015), Objaverse-XL Deitke et al. (2023a)), subjected to the same rigorous filtering pipeline described in Section A. This experiment serves to answer a critical question: can a robust, scalable methodology achieve state-of-the-art results even when limited to open-source data?
The results, shown in Figure 15, are twofold and unequivocal. First, QuadGPT-OS still significantly outperforms all prior autoregressive baselines. This confirms that our core contributions—the scalable architecture, curriculum learning strategy, and tDPO refinement—are the primary drivers of our model’s high performance. It proves that our method is not merely a product of better data, but a fundamentally more solid approach to the problem.
Second, our full model, trained on the complete 1.3 million mesh dataset, demonstrates a further substantial improvement in quality over QuadGPT-OS. This highlights a key insight: while a superior methodology can establish a new state of the art on public data, the full potential of a truly scalable approach is unlocked through careful, large-scale data curation. Our work demonstrates the synergistic impact of advancing both the model and the data to push the frontier of production-ready 3D generation.
Figure 15: Qualitative results of our data ablation study. “QuadGPT-OS” is trained exclusively on filtered public datasets, while “QuadGPT” uses our complete curated dataset. “QuadGPT-OS” already demonstrates a significant improvement in topological coherence and geometric detail over prior baselines. The full model shows further refinement, particularly in achieving cleaner, more professional edge flow, validating the scalability of our approach with high-quality data.